Find the minimum value for the function: y = x^2 + 9x + 18
Answer by lwsshak3(11628)

Find the minimum value for the function:
y = x^2 + 9x + 18
This is an equation of a parabola that opens upwards (function has a minimum)
Its standard form: y=(x-h)^2+k, (h,k)=(x,y) coordinates of vertex, k=minimum value
complete the square
y=(x^2+9x+20.25)+18-20.25
y=(x+9/2)^2-2.25
minimum value=-2.25